PLEASE SHOW STEP BY STEP THANK YOU

F Corporation purchased equipment for $450,000. Residual value at the end of the estimated 6 years or 150,000 hours service life is expected to be $60,000.

Using Straight-line depreciation, determine:

  1. Book value at the end of year 3      $_____

Using  180% declining balance depreciation, determine:

  1. Depreciation expense for year 2     $_____
  2. Depreciation expense for year 3     $_____

Year

Hours used

1

25,000

2

24,900

3

25,100

4

25,400

5

25,100

6

24,500

Using Activity based depreciation, determine:

  1. Depreciation expense for year 3    $____
  2. Book value at the end of year 3      $____
Answer #1

1) Using straight line

Book value at the end of year 3 = 450000-(450000-60000/6*3) = 255000

2) Using double decline

Double decline rate = 100/6*1.8 = 30%

Depreciation expense for year 2 = 450000*70%*30% = 94500

Depreciation expense for year 3 = 450000*70%*70%*30% = 66150

3) Using activity based

Depreciation expense for year 3 = (450000-60000/150000)*25100 = 65260

Book value at the end of year 3 = 450000-(75000*2.6) = 255000